# Stochastic Volatility in Mean: Efficient Analysis by a Generalized Mixture Sampler

### v1

@@ Abstract
In this paper we consider the simulation-based Bayesian analysis of stochastic volatility in mean (SVM) models. Extending the highly efficient Markov chain Monte Carlo mixture sampler for the SV model proposed in Kim et al. (1998) and Omori et al. (2007), we develop an accurate approximation of the non-central chi-squared distribution as a mixture of thirty normal distributions. Under this mixture representation, we sample the parameters and latent volatilities in one block. We also detail a correction of the small approximation error by using additional Metropolis-Hastings steps. The proposed method is extended to the SVM model with leverage. The methodology and models are applied to excess holding yields in empirical studies, and the SVM model with leverage is shown to outperform competing volatility models based on marginal likelihoods.
@@
Keywords: Excess Holding Yield; GARCH in Mean; EGARCH-in Mean; Markov chain Monte Carlo; Mixture Sampler; Risk Premium; Stochastic Volatility in Mean
@@ 1 Introduction
In the existing literature, the SVM model without leverage has been estimated by the multi-move (block) sampler (Shephard and Pitt (1997) and Omori and Watanabe (2008)), for example, Abanto-Valle et al. (2011), Abanto-Valle et al. (2012), and Leão et al. (2017), and by other approaches, for example, Chan (2017), Abanto-Valle et al. (2021), and Abanto-Valle et al. (2023). There is no known mixture sampler approach for SVM models with leverage. In this paper, we develop efficient MCMC based algorithms for SVM models, with and without leverage, that are based on accurate representations of these models in terms of mixtures of conditionally Gaussian linear state-space models, just as in the approach of Kim et al. (1998). However, due to the $\beta\exp(h_{t}/2)$ term in the mean equation, instead of characterizing the distribution of a central chi-squared distribution with one degrees of freedom in terms of mixtures of normal distributions, a mixture representation to the distribution of a non-central chi-squared distribution with one degrees of freedom is needed, in which the non-centrality parameter is $\beta^{2}$. We show that the latter distribution has an infinite series expansion. Our estimation approach uses a truncated version of this series expansion to develop a highly efficient fitting algorithm. It can be viewed as a generalized mixture sampler. Furthermore, the small approximation error due to the truncation can be corrected by a data augmentation method by incorporating a pseudo-target probability density whose marginal probability density is the exact conditional posterior density.
We apply the proposed method to excess holding yields data, where we also consider alternative ways of introducing the log-variance in the mean function. We show that our proposed SVM model outperforms other competing volatility models such as the basic SV model, alternative SVM models, GARCH models, GARCH in mean models, EGARCH models and EGARCH-in mean models based on marginal likelihoods.
The rest of this paper is organized as follows. Section 2 introduces the SVM model and describes the novel mixture sampler as an efficient sampling method for such models. Section 3 illustrates the performance of this sampling method using the simulated data for several cases. Section 4 further extends the SVM model to incorporate the leverage effect. Finally, in Section 5, we apply our proposed SVM model to financial data and perform a comprehensive evaluation of the model. Conclusion and remarks are given in Section 6.
@@ 2.1 SVM model
We define the stochastic volatility in mean (SVM) model as follows.
@@
where $N(m,S)$ denotes normal distribution with mean vector $m$ and covariance matrix $S$, $\theta=(\mu,\phi,\sigma^{2},\beta)$ is a model parameter vector of interest, and $h=(h_{1},\dots,h_{n})^{\prime}$ is the logarithm of the latent volatility vector. We use $\beta\exp(h_{t}/2)$ rather than $\beta\exp(h_{t})$ in the mean equation since $\beta$ in our specification is not affected by the change of measurement units. For $\beta\neq 0$, we denote it as the stochastic volatility in mean (SVM) model. The standard stochastic volatility (SV) model is obtained as a special case with $\beta=0$. For $\theta$, we assume the prior distributions
@@ 2.2 Transformation of the measurement equation
$f(\epsilon_{t}^{*})=\frac{1}{\sqrt{2\pi}}\left(\frac{\epsilon_{t}^{*}-\exp(\epsilon_{t}^{*})}{2}\right),\quad-\infty<\epsilon_{t}^{*}<\infty,$
@@
This expression implies that the probability density function of $\log\chi_{1}^{2}(\lambda)$ is approximated by the mixture of $K(J+1)$ normal distributions. Especially, when $\lambda=0$ and $J=0$, the approximation (9) reduces to (8). The extent of the impact on the approximation is based on the value $w_{i,j}$ which includes $\lambda=\beta^{2}$. The coefficient $\beta$ of volatility $\exp(h_{t}/2)$ is often estimated to be $\hat{\beta}=0.1\sim 0.2$ in past empirical studies for GARCH-M (generalized autoregressive conditional heteroscedasticity in mean) models. When $\lambda=\beta^{2}$ is less than 1.0, $J=1$ or $2$ makes $\sum_{i=1}^{K}\sum_{j=0}^{J}w_{i,j}$ more than 0.9. For $J=2$, Figures 1 and 2 show the true and approximate densities of $\log\chi_{1}^{2}(\beta^{2})$ for $\beta=0.3,0.5$ and $0.7$ (equivalently, $\beta=-0.3,-0.5$ and $-0.7$) and the difference between the two densities, respectively. Since these differences are quite small and the approximation almost overlaps the true probability density of $\log\chi_{1}^{2}(\beta^{2})$ even for $\beta=0.7$, we employ $J=2$ in this paper. That is, we approximate the probability density of $\epsilon_{t}^{*}|\beta\sim\log\chi_{1}^{2}(\beta^{2})$ by
@@ 3.1 MCMC algorithm
Algorithm 1. Let us denote $\theta=(\alpha,\beta)$ where $\alpha=(\mu,\phi,\sigma^{2})$. The Markov chain Monte Carlo simulation is implemented in four blocks:
@@
As discussed in the previous section, we sample $h$ using the mixture sampler using the mixture of normal distributions. Since our approximation is highly accurate, we incorporate this mixture approximation in Step 3 without correcting the approximation error as in Step 2 of Algorithm 1 in Chib et al. (2002). The additional step to correct for the approximation error is described in Algorithm 2 in the Appendix A. Let $f_{N}(\cdot|m,s^{2})$ denote the probability density of $N(m,s^{2})$, and let $\pi(\alpha)$ denote the prior density of $\alpha$. Define our target density in Step 3 as
@@ 3.2 Associated particle filter
is a consistent estimate of the conditional log-likelihood and can be used as an input in the calculation of the marginal likelihood by the method of Chib (1995).
@@ 4 Illustrative numerical examples
$\mu\sim N(0,1000^{2}),\quad\frac{\phi+1}{2}\sim Beta(1,1),$
@@
    Case $\beta=0.3$. The acceptance rates of the MH algorithms for $\alpha$ is 73.6%. The sample paths are given in Figure 3, and the MCMC chain mixes quite well.
@@
    Case $\beta=0.5$. The acceptance rates of the MH algorithms for $\alpha$ is 73.6%. The plot of the sample paths and log volatilities are similar to those in (i) and hence omitted to save space. Table 3 shows the posterior means, 95% credible intervals and inefficiency factors. The estimated parameters are close to true values, and inefficiency factors (IF) are overall small as in (i). The IF’s for $h_{250}$ and $h_{750}$ are sufficiently small, and indicates that the algorithm is still highly efficient.
@@
    Case $\beta=0.7$. The acceptance rates of the MH algorithms for $\alpha$ is 73.5%. The convergence seems to become slightly slower, but the chain mixes well. The plot of the sample paths and log volatilities are similar to those in (i) and hence omitted to save space. Table 9 shows the posterior means, 95% credible intervals and inefficiency factors. The estimated parameters are close to true values, and inefficiency factors (IF) are overall relatively small. The IF’s for $h_{250}$ and $h_{750}$ are small, and indicates that the algorithm is still works well.
@@ 6.1 Data
This section applies the SVM model with leverage and several alternative models to three excess holding yields data. Further, we conduct a comprehensive model comparison including the GARCH, GARCH in mean, EGARCH and EGARCH in mean models. The descriptions of the data (labeled as TB, DGS and AAA) are given below [^1] The data are obtained from the website of Federal Reserve Bank of St. Louis..
@@
    AAA: the excess holding yield using 3 months and 20 years Moody’s Seasoned AAA corporate bond yield with 359 observations from the fourth quarter of 1933 to the second of 2023. As discussed in Engle et al. (1987), assuming that the bonds are effectively infinitely lived, we define
    $y_{t}=\left\{\frac{R_{t}}{100}-\frac{r_{t}}{100}-1+\frac{R_{t}}{R_{t+1}}\right\}\times 100,$
    at annual rate where $R_{t}$ is the Moody’s seasoned AAA corporate bond yield (percent, monthly, not seasonally adjusted) and $r_{t}$ is the 3-month treasury bill secondary market rate (percent, discount basis, monthly, not seasonally adjusted), measured at the beginning of the quarter.
@@
Figure 8: Time series plot. AAA data: 1933Q4–2023Q2.

### v2

@@ Abstract
In this paper we consider the simulation-based Bayesian analysis of stochastic volatility in mean (SVM) models. Extending the highly efficient Markov chain Monte Carlo mixture sampler for the SV model proposed in Kim et al. (1998) and Omori et al. (2007), we develop an accurate approximation of the non-central chi-squared distribution as a mixture of thirty normal distributions. Under this mixture representation, we sample the parameters and latent volatilities in one block. We also detail a correction of the small approximation error by using additional Metropolis-Hastings steps. The proposed method is extended to the SVM model with leverage. The methodology and models are applied to excess holding yields and S&P500 returns in empirical studies, and the SVM models are shown to outperform other volatility models based on marginal likelihoods.
@@
Keywords: Excess Holding Yield; Markov chain Monte Carlo; Mixture Sampler; Risk Premium; Stochastic Volatility in Mean
@@ 1 Introduction
In the existing literature, the SVM model without leverage has been estimated by the multi-move (block) sampler (Shephard and Pitt (1997) and Omori and Watanabe (2008)), for example, Abanto-Valle et al. (2011), Abanto-Valle et al. (2012), and Leão et al. (2017), and by other similar approaches, for example, Chan (2017), Abanto-Valle et al. (2021), and Abanto-Valle et al. (2023). There is no known mixture sampler approach for SVM models with leverage. In this paper, we develop efficient MCMC based algorithms for SVM models, with and without leverage, that are based on accurate representations of these models in terms of mixtures of conditionally Gaussian linear state-space models, just as in the approach of Kim et al. (1998). However, due to the $\beta\exp(h_{t}/2)$ term in the mean equation, instead of characterizing the distribution of a central chi-squared distribution with one degrees of freedom in terms of mixtures of normal distributions, a mixture representation to the distribution of a non-central chi-squared distribution with one degrees of freedom is needed, in which the non-centrality parameter is $\beta^{2}$. We show that the latter distribution has an infinite series expansion. Our estimation approach uses a truncated version of this series expansion to develop a highly efficient fitting algorithm. It can be viewed as a generalized mixture sampler. Furthermore, the small approximation error due to the truncation can be corrected by a data augmentation method by incorporating a pseudo-target probability density whose marginal probability density is the exact conditional posterior density.
We apply our proposed method to excess holding yields and S&P500 returns data, and show that SVM models outperform other SV models based on marginal likelihoods. The rest of this paper is organized as follows. Section 2 introduces the SVM model and describes the novel mixture sampler as an efficient sampling method for such models. The MCMC simulation and the particle filter are described in Section 3. Section 4 illustrates the performance of this sampling method using the simulated data for several cases. Section 5 further extends the SVM model to incorporate the leverage effect. Finally, in Section 6, we apply our proposed SVM model to financial data and perform a model comparison. Conclusion and remarks are given in Section 7.
@@ 2.1 SVM model
We define the stochastic volatility in mean (SVM) model as follows:
@@
where $N(m,S)$ denotes normal distribution with mean vector $m$ and covariance matrix $S$, $\theta=(\mu,\phi,\sigma^{2},\beta)$ is a model parameter vector of interest, and $h=(h_{1},\dots,h_{n})^{\prime}$ is the logarithm of the latent volatility vector. We use the standard deviation $\exp(h_{t}/2)$ in the mean equation, rather than the variance $\exp(h_{t})$, to match the units of the outcome variable. For $\beta\neq 0$, we denote it as the stochastic volatility in mean (SVM) model. The standard stochastic volatility (SV) model is obtained as a special case with $\beta=0$. For $\theta$, we assume the prior distribution
@@ 2.2 Transformation of the measurement equation
$f(\epsilon_{t}^{*})=\frac{1}{\sqrt{2\pi}}\exp\left(\frac{\epsilon_{t}^{*}-\exp(\epsilon_{t}^{*})}{2}\right),\quad-\infty<\epsilon_{t}^{*}<\infty,$
@@
This expression implies that the probability density function of $\log\chi_{1}^{2}(\lambda)$ is approximated by the mixture of $K(J+1)$ normal distributions. Especially, when $\lambda=0$ and $J=0$, the approximation (9) reduces to (8). The extent of the impact on the approximation is based on the value $w_{i,j}$ which includes $\lambda=\beta^{2}$. The coefficient $\beta$ of volatility $\exp(h_{t}/2)$ is estimated to be less than one in past empirical studies. When $\lambda=\beta^{2}$ is less than 1.0, $J=1$ or $2$ makes $\sum_{i=1}^{K}\sum_{j=0}^{J}w_{i,j}$ more than 0.9. For $J=2$, Figures 1 and 2 show the true and approximate densities of $\log\chi_{1}^{2}(\beta^{2})$ for $\beta=0.3,0.5$ and $0.7$ (equivalently, $\beta=-0.3,-0.5$ and $-0.7$) and the difference between the two densities, respectively. Since these differences are quite small and the approximation almost overlaps the true probability density of $\log\chi_{1}^{2}(\beta^{2})$ even for $\beta=0.7$, we employ $J=2$ in this paper. That is, we approximate the probability density of $\epsilon_{t}^{*}|\beta\sim\log\chi_{1}^{2}(\beta^{2})$ by
@@ 3.1 MCMC algorithm
Algorithm 1 (Generalized mixture sampler, GMS). Let us denote $\theta=(\alpha,\beta)$ where $\alpha=(\mu,\phi,\sigma^{2})$. The Markov chain Monte Carlo simulation is implemented in four blocks:
@@
As discussed in the previous section, we sample $h$ using the mixture sampler using the mixture of normal distributions. Since our approximation is highly accurate, we can use this mixture approximation directly in Step 3, as in Step 2 of Algorithm 1 in Chib et al. (2002). However, one can remove the small approximation error with an additional MH step, as detailed in Algorithm 2, GMS with MH algorithm (GMH), given in the Appendix A, but due to the fact that the tailored mixture very closely fits the non-central chi-squared distribution, this additional step would be rarely necessary.
Let $f_{N}(\cdot|m,s^{2})$ denote the probability density of $N(m,s^{2})$, and let $\pi(\alpha)$ denote the prior density of $\alpha$. Define our target density in Step 3 as
@@ 3.2 Associated particle filter
is a consistent estimate of the conditional log-likelihood and can be used as an input in the calculation of the marginal likelihood by the method of Chib (1995), as extended to M-H sampler output by Chib and Jeliazkov (2001).
@@ 4 Illustrative numerical examples
$\mu\sim N(0,3^{2}),\quad\frac{\phi+1}{2}\sim Beta(1,1),$
@@
    Case $\beta=0.3$. The acceptance rates of the MH algorithms for $\alpha$ is 72.8%. The sample paths are given in Figure 3, and the MCMC chain mixes quite well.
@@
    Case $\beta=0.5$. The acceptance rates of the MH algorithms for $\alpha$ is 73.2%. The plot of the sample paths and log volatilities are similar to those in (i) and hence omitted to save space. Table 3 shows the posterior means, 95% credible intervals and inefficiency factors. The estimated parameters are close to true values, and inefficiency factors (IF) are overall small as in (i). The IF’s for $h_{250}$ and $h_{750}$ are sufficiently small, and indicates that the algorithm is still highly efficient.
@@
    Case $\beta=0.7$. The acceptance rates of the MH algorithms for $\alpha$ is 75.2%. The convergence seems to become slightly slower, but the chain mixes well. The plot of the sample paths and log volatilities are similar to those in (i) and hence omitted to save space. Table 11 shows the posterior means, 95% credible intervals and inefficiency factors. The estimated parameters are close to true values, and inefficiency factors (IF) are overall relatively small. The IF’s for $h_{250}$ and $h_{750}$ are small, and indicates that the algorithm is still works well.
@@
Comparison of sampling efficiencies
Next, we compare the sampling efficiency of our proposed method with the method in Chan (2017), which was shown to be more efficient than the more complex methods of McCausland (2012) and Andrieu et al. (2010). Chan’s method needs to implement the accept-reject Metropolis-Hastings (ARMH) algorithm (Chib and Greenberg, 1995) (the usual MH algorithm could be used, but then the rejection rate becomes much higher), using a proposal distribution based on a multivariate normal approximation of the full conditional distribution, obtained by a second-order Taylor expansion.
Table 5 shows the IFs for selected values ($h_{t}$, $t=100,200,\ldots,1000$), the mean ($\overline{h}$) and the median ($h_{med}$) of $h_{t}$ for $\beta=0.3,0.5$ and $0.7$. Among the three algorithms, the generalized mixture sampler (Algorithm 1, denoted by GMS) is the most efficient with IFs less than 10. These are substantially smaller than those of the CHN method. The IFs for Algorithm 2, the generalized mixture sampler with an additional Metropolis-Hastings step (denoted as GMH), are somewhat larger, but still smaller than those from Chan’s method (denoted by CHN). We note that the IFs of GMH become larger as the absolute value of $\beta$ increases. However, as shown in the results of the simulation experiment above and in Appendix A, the resulting posterior estimates by GMS and GMH are almost the same. The takeaway from this experiment is the striking efficiency of the GMS method.
In addition, we compare the computational times required in each experiment for the three algorithms. As shown in Table 6, our proposed methods (GMS and GMH) are much faster than Chan’s method (CHN). This is mainly because (i) GMS and GMH sample $h$ using the highly efficient and fast simulation smoother, whereas the CHN method is based on an expensive tailoring step that entails the inversion of a high dimensional covariance matrix in each MCMC iteration; (ii) the AR step of the ARMH algorithm degrades with many rejections when $h$ is high-dimensional and $\beta$ is large; and (iii) the search for the mode in the tailoring step by an iterative optimization methods consumes considerable time for longer time series (i.e., as the dimension of $h$ increases). Our proposed methods encounters none of these difficulties.
$\beta=0.3$ $\beta=0.5$ $\beta=0.7$
$h_{t}$ GMS GMH CHN GMS GMH CHN GMS GMH CHN
$h_{100}$ 6 11 32 2 26 18 2 94 87
$h_{200}$ 4 32 50 5 21 40 6 85 89
$h_{300}$ 2 12 22 7 31 43 3 116 73
$h_{400}$ 4 23 67 3 41 63 4 82 65
$h_{500}$ 8 10 43 6 32 60 6 88 209
$h_{600}$ 1 17 41 2 35 48 2 38 40
$h_{700}$ 5 18 50 5 26 15 4 51 64
$h_{800}$ 4 12 38 6 71 14 3 56 103
$h_{900}$ 3 11 27 7 48 30 4 25 26
$h_{1000}$ 2 6 23 3 35 42 5 114 99
$\overline{h}$ 8 28 182 9 68 105 9 135 165
$h_{med}$ 7 15 74 7 25 35 4 62 70
Table 5: Inefficiency factors of selected $h_{t}$’s with their mean ($\overline{h}$) and median ($h_{med}$) in our simulation studies. Generalized mixture sampler (GMS), GMS with MH step for approximation correction (GMH), and Chan’s method (Chan (2017)) (CHN).
$\beta=0.3$ $\beta=0.5$ $\beta=0.7$
GMS GMH CHN GMS GMH CHN GMS GMH CHN
Time 1,792 1,910 133,884 1,701 1,935 134,127 1,700 1,948 126,168
Table 6: Computational time (seconds) for our simulation studies. We draw 50,000 MCMC samples after discarding initial 10,000 samples as burn-in period.
@@ 5 Extension to SVM model with Leverage (SVML model)
Extension to the multivariate SVM model
Furthermore, we will give a couple of examples to illustrate how to extend our proposed SVM model to the multivariate models. As a first example, consider the factor multivariate stochastic volatility (MSV) model proposed by Chib et al. (2006) (see Ishihara and Omori (2017) for the model with leverage). Let $\bm{y}_{t}=(y_{1t},\ldots,y_{pt})^{\prime}$ and $\bm{f}_{t}=(f_{1t},\ldots,f_{qt})^{\prime}$ denote the dependent and factor variables ($q<p$) respectively. The basic factor MSV model is given by
$\bm{y}_{t}$ $=$ $\mathbf{B}\bm{f}_{t}+\mathbf{V}_{1t}^{1/2}\bm{\epsilon}_{1t},\quad\bm{\epsilon}_{1t}\sim N(\bm{0},\mathbf{I}_{p}),$
$\bm{f}_{t}$ $=$ $\mathbf{V}_{2t}^{1/2}\bm{\epsilon}_{2t},\quad\bm{\epsilon}_{2t}\sim N(\bm{0},\mathbf{I}_{q}),$
$\bm{h}_{t+1}$ $=$ $\bm{\mu}+\mathbf{\Phi}(\bm{h}_{t-1}-\bm{\mu})+\bm{\eta}_{t},\quad\bm{\eta}_{t}\sim N(\bm{0},\mathbf{\Sigma}),$
$\mathbf{V}_{1t}=\mbox{diag}(\exp(h_{1t}),\ldots,\exp(h_{pt})),\quad\mathbf{V}_{2t}=\mbox{diag}(\exp(h_{p+1,t}),\ldots,\exp(h_{p+q,t})),$
$\mathbf{\Phi}=\mbox{diag}(\phi_{1},\ldots,\phi_{p}),\quad\mathbf{\Sigma}=\mbox{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2},\sigma_{p+1}^{2},\ldots\sigma_{p+q}^{2}),$
where $\mathbf{B}$ is a $p\times q$ factor loading matrix subject to constraints ($b_{ij}=0$ for $j>i$ and $b_{ii}=1$ for $i\leq q$). To incorporate the SVM in the MSV model, we replace the factor equation by
$\bm{f}_{t}$ $=$ $\mathbf{V}_{2t}^{1/2}\bm{\beta}+\mathbf{V}_{2t}^{1/2}\bm{\epsilon}_{2t},$
where $\bm{\beta}=(\beta_{1},\ldots,\beta_{q})^{\prime}$ is a $q\times 1$ vector of weights.
As a second example, we may consider the MSV model with the SVM that is common within a group. Let $\bm{y}_{t}=(y_{1t},\ldots,y_{pt})^{\prime}$ and consider
$\bm{y}_{t}$ $=$ $\exp(\tilde{h}_{t}/2)\bm{\beta}+\exp(\tilde{h}_{t}/2)\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim N(\bm{0},\mathbf{\Sigma}),$
$\tilde{h}_{t+1}$ $=$ $\phi\tilde{h}_{t}+\eta_{t},\quad\eta_{t}\sim N(0,\sigma^{2}),\quad\mathbf{\Sigma}=\mbox{diag}(\sigma_{1}^{2},\ldots,\sigma_{p}^{2}),$
where $\bm{\beta}=(\beta_{1},\ldots,\beta_{p})^{\prime}$. To linearize the measurement equation with respect to $\tilde{h}_{t}$, we define $\bm{y}^{*}=(y_{1t}^{*},\ldots,y_{pt}^{*})^{\prime}$ and $\bm{\epsilon}_{t}^{*}=(\epsilon_{1t}^{*},\ldots,\epsilon_{pt}^{*})^{\prime}$ where $y_{it}^{*}=\log y_{it}^{2}$ and $\epsilon_{it}^{*}=\log(\beta_{i}/\sigma_{i}+\epsilon_{it}/\sigma_{i})^{2}$. Thus the transformed measurement equation is
$\bm{y}_{t}^{*}=\bm{\mu}^{*}+\tilde{h}_{t}\bm{1}_{p}+\bm{\epsilon}_{t}^{*},\quad\bm{\mu}^{*}=(\log\sigma_{1}^{2},\ldots,\log\sigma_{p}^{2})^{\prime},\quad\bm{1}_{p}=(1,\ldots,1)^{\prime},$
Noting that $\epsilon_{it}^{*}\sim\log\chi_{1}^{2}(\beta_{i}^{2}/\sigma_{i}^{2})$, we can approximate the noncentral chisquare distribution by the mixture of normal distributions as we have described.
Remark. Some previous studies considered alternative mean specifications of $y_{t}$ using $\exp(h_{t})$ or $h_{t}$, as in Engle et al. (1987) for ARCH-M model. We note that our proposed sampler could be utilized to generate $h_{t}^{\prime}s$ efficiently, but details are left for our future work.
@@ 6.1 Data
This section applies the SVM model with leverage and several alternative models to three excess holding yields data. The descriptions of the data (labeled as TB, DGS and S&P500) are given below [^1] The data are obtained from the website of Federal Reserve Bank of St. Louis..
@@
    S&P500: the excess return using S&P500 index daily return and federal funds rate with 1008 observations from July 1st of 2019 to June 30 of 2023. It is defined as $y_{t}=R_{t}-r_{t}$ at daily rate where $R_{t}$ and $r_{t}$ are the daily log return of S&P500 (in percent), and federal funds effective rate (percent, daily, not seasonally adjusted) divided by 360.
@@
Figure 8: Time series plot. S&P500 data: 2019/7/1–2023/6/30.